# Node-opening construction of maxfaces

### v2

@@ Abstract.
The node-opening technique developed by Traizet has been very useful in constructing minimal surfaces. In this paper, we use the technique to construct families of maximal immersions in Lorentz space that are embedded outside a compact set. Each family depends on a real parameter $t$. The surfaces look like spacelike planes connected by small necks that shrink to singular points as $t\to 0$. The limit positions of the necks must satisfy a balance condition, which turns out to be exactly the same for maxfaces and for minimal surfaces. By simply comparing notes, we obtain a rich variety of new maxfaces with high genus and arbitrarily many space-like ends, including the Lorentzian Costa–Hoffman–Meeks surfaces, thereby remove the scarcity of examples of maxfaces. Moreover, a main task of this paper is to analyse the singularities, an interesting feature that distinguishes maxfaces from minimal surfaces. For sufficiently small non-zero $t$, the singular set consists of curves in the waist of every neck. In generic and some symmetric cases, we are able to prove that all but finitely many singularities are cuspidal edges, and the non-cuspidal singularities are swallowtails.
@@ 1. Introduction
Maximal surfaces are zero mean curvature immersions in the Lorentz Minkowski space ${\mathbb{E}}_{1}^{3}$. These surfaces emerge as solutions to the variational problem of locally maximizing the area among spacelike surfaces. They share several similarities with minimal surfaces in ${\mathbb{R}}^{3}$. For instance, both are critical points of the area functional and admit Weierstrass-Enneper representations. However, while there are rich examples of complete minimal surfaces, the only complete maximal immersion is the plane [umehara2006].
@@
Now that singularities are allowed, there are many examples of complete maxfaces, such as the Lorentzian Catenoid and Kim-Yang Toroidal maxface [kim2006]. In 2006, Kim and Yang [kim2006] constructed complete maximal maps of genus $k\geq 1$. When $k=1$, it is a complete embedded (in a wider sense) maxface known as Kim-Yang toroidal maxface. When $k>1$, they are not maxfaces. In [fujimori2009], the authors constructed a family of complete maxfaces $f_{k}$ for $k\geq 1$ with two ends; but when $k>1$, these may not be embedded (in a wider sense). Moreover, in 2016, Fujimori, Mohamed, and Pember [fujimori2016h] constructed maxfaces of any odd genus $g$ with two complete ends (if $g=1$, the ends are embedded) and maxfaces of genus $g=1$ and three complete embedded ends.
@@
There is a scarcity of examples of complete maxfaces [fujimori2009]. All higher-genus maxfaces in the literature, to our best knowledge, have only two or three ends, usually of the catenoid type. Very few known higher-genus maxfaces are embedded (in a wider sense). Open problems in [fujimori2009] expressed hope for a large collection of examples of complete maxfaces that are embedded (in a wider sense) with higher genus and many ends.
@@
In this paper, we construct a rich variety of complete maxfaces that are embedded (in a wider sense) of high genus and arbitrary number of spacelike ends. Each of these families depends on a sufficiently small non-zero real parameter $t$.
@@
Our primary tool of construction will be the node opening technique, a Weierstrass gluing method developed by Traizet [traizet2002e]. The approach starts from Weierstrass data defined on a Riemann surface with nodes, then obtains Riemann surfaces by opening the nodes into necks and, at the same time, extends the Weierstrass data to Riemann surfaces using the Implicit Function Theorem. The node-opening technique has been very successful in constructing a rich variety of minimal surfaces [traizet2002e]. However, to the best of our knowledge, the current paper marks the first application of the technique to surfaces in the Lorentz-Minkowski space.
@@
The Weierstrass gluing method has several advantages over other methods. On the one hand, in the existing literature on maxfaces, authors often need to assume symmetries to make the construction possible, hence only produce examples restricted to symmetries. The gluing technique has been a very powerful tool to break symmetries in minimal surfaces; in some sense, the technique was developed for this purpose [traizet2002e]. We will see later that it is equally powerful in breaking symmetries for maximal surfaces, hence ideal for removing the scarcity of examples. On the other hand, while the PDE gluing method is also popular for constructing minimal surfaces, the existence of singularities makes it difficult to apply on maxfaces. More specifically, one needs to identify (glue) two curves in the process, but the analysis would be difficult if the curves contain singularities. In the Weierstrass gluing process, we instead identify (glue) two annuli; hence, we can bypass the singularities.
@@
The node-opening construction for maxfaces turns out to be very similar to that for minimal surfaces, hence will only be sketched in Section 4. This similarity implies that, for any minimal surface constructed by opening nodes, there is a corresponding maxface also constructed by opening nodes. This correspondence between maximal and minimal immersions is different from the usual correspondence through Weierstrass data [umehara2006]. Then in Section 3, by simply comparing notes, we obtain a rich variety of new maxfaces with high genus and arbitrarily many space-like ends, thereby remove the scarcity of examples of maxfaces. Among the examples are the Lorentzian Costa and Costa–Hoffman–Meeks (CHM) surfaces and their generalizations with arbitrarily many ends, providing positive answers to the open problems in [fujimori2009]. To the best of our knowledges, these are the first time that Lorentzian analogues of CHM surfaces were constructed.
@@
The singularity structure is an interesting feature that is special for maxfaces. Complete non-planar maxfaces always appear with singularities, such as cuspidal edges, swallowtails, cuspidal crosscaps, and cone-like singularities, to name a few. We refer the readers to [umehara2006, sai2022, kumar2020] to explore various singularities on maxfaces. In Section 5, we will analyse the types of singularities on the maxfaces that we construct using their Weierstrass data.
Components of the singular set are waists around the necks. We first prove in Theorem 5.1 that, around a specific neck and for sufficiently small non-zero $t$, either the singular set is mapped to a single point (cone-like singularity), or all but finitely many singular points are cuspidal edges. Moreover, the finitely many non-cuspidal singularities are generalized $A_{k}$ singularities, their positions on the waist depend analytically on $t$, and their types do not vary for sufficiently small $t$. Then, in Proposition 5.2, we show that, generically, there are four swallowtail singularities around a neck. The non-generic cases are generally hard to study. But in Section 5.5, we managed to analyse the singularities in the presence of rotational and reflectional symmetries. In Section 3, we will use the Lorentzian Costa and Costa–Hoffman–Meeks surfaces to examplify our results on singularities.
@@ 2. Main results
We now state our main result.
@@
To state our result on singularities, we need to define
@@
Our results about singularities are summarized below.
Theorem (Informal).
@@
    Assume that the configuration has rotational symmetry of order $r>1$s around a neck and $R^{(r-1)}_{l,k}\neq 0$, then for sufficiently small non-zero $t$, there are $2r$ swallowtail singularities around the neck, and they tend to be evenly distributed as $t\to 0$.
@@ 3.2. Singularities on Lorentzian Costa and CHM surfaces
4. Sketched construction
4.1. The Weierstrass data
@@
4.1.1. The Riemann Surface
@@
4.1.2. Gauss map and local coordinates
@@ 4.1.3. The height differential
$\Omega_{l}:=\{z\in{\mathbb{C}}_{l}:|v_{l,i}|<\varepsilon\quad\forall 1\leq i\leq n_{l}\quad\text{and}\quad|w_{l-1,j}|<\varepsilon\quad\forall 1\leq j\leq n_{l-1}\},$
@@
So, it suffices to prescribe the residue $-R_{k}$ and the periods around $\gamma_{l,k}$ for $1<k\leq n_{l}$.
@@
4.2. Using the Implicit Function Theorem
@@
Proposition 4.2 (Divisor condition).
@@
Proposition 4.3 (Vertical periods).
@@
Proposition 4.4 (Horizontal periods).
Given a balanced and rigid configuration $(p,Q)$ such that the map $Q\to W$ has rank $1$. Assume that $(\alpha,\beta,r)$ are given by previous propositions. For $\tau$ in a neighborhood of 00, there exists unique values for $a$, $b$, and $R$, depending smoothly on $\tau$, such that $\sum R_{l}=0$ and the horizontal period condition (4.2) are satisfied over the curves $\Gamma_{l,k}$ and $\gamma_{l,k}$, $1\leq l<L$ and $1<k\leq n_{l}$. Moreover, at $\tau=0$, up to a translation in ${\mathbb{C}}_{l}$, we have $a_{l,k}=\overline{b_{l,k}}=\overline{p_{l,k}}$ if $l$ is odd, $=p_{l,k}$ if $l$ is even, and $R_{l}=Q_{l}$.
@@
5. Singularities
with $1\leq l\leq N-1,1\leq k\leq n_{l}$. In this section, we aim to analyze the nature of these singularities.
5.1. The governing function
We need to study the function
where $f_{t}:=dh_{t}/dv$.
On the other hand, it was shown in [umehara2006] that
In particular, $p=te^{i\theta}$ is a cuspidal singularity whenever
and $p$ is a swallowtail singularity whenever
5.2. Nondegeneracy
5.3. Non-cuspidal singularities
We have proved the following
Theorem 5.1.
5.4. Swallowtails
We first remove the trivial solutions $t=0$ by considering the function
In the Appendix, we compute that the partial derivatives
where
In particular,
Note that, at $t=0$, we have
Then, we must look at the next derivative, namely
Proposition 5.2.
5.5. Symmetries
Proposition 5.3.
Proof.
Proposition 5.4.
Proof.
Remark 5.5.
Remark 5.6.
Proposition 5.7.
Proof.
Appendix A Derivatives of ${\mathcal{A}}$
The height differential $dh_{t}$ for the maxface, as defined in Section 4.1.3, has the following Laurent expansion in the annulus $t^{2}/\varepsilon<v<\varepsilon$:
where
Moreover, for each $n\in{\mathbb{Z}}$,
$a_{n}=\frac{1}{2\pi i}\int_{|v|=\varepsilon}\frac{dh_{t}}{v^{n+1}}=\frac{-1}{2\pi i}\int_{|w|=\varepsilon}\frac{dh_{t}w^{n+1}}{t^{2n+2}}.$
Therefore, we have
(A.1) $\frac{\partial^{m}a_{n}}{\partial t^{m}}=\frac{-1}{2\pi i}\sum_{j=0}^{m}\int_{|w|=\varepsilon}w^{n+1}\binom{m}{j}\frac{\partial^{m-j}dh_{t}}{\partial t^{m-j}}(-2n-2)_{j}\frac{1}{t^{2n+2+j}}.$
We now prove Equation (5.6) for the partial derivatives of ${\mathcal{A}}$ over $t$. We repeat the formula below
Proof.
Recall the expression (5.1) of ${\mathcal{A}}$
(A.2) ${\mathcal{A}}(t,\theta)=te^{i\theta}\frac{dh_{t}}{dv}(te^{i\theta})=\sum_{n\in{\mathbb{Z}}}a_{n}(t)t^{n+1}e^{i(n+1)\theta}=\sum_{n\in{\mathbb{Z}}}b_{n}(t)t^{n+1}e^{-i(n+1)\theta},$
which can be expanded in terms of $t$ as follows
(A.3) ${\mathcal{A}}(t,\theta)=\sum_{m\geq 0}\frac{1}{m!}\frac{\partial^{m}{\mathcal{A}}}{\partial t^{m}}(0,\theta)t^{m}.$
From (A.2), we have
$n>m-1$:
$n=m-1$:
In this case, as $t\to 0$, we have
since derivatives of $a_{n}$ are bounded and, for $j<m$, $t^{m-j}\to 0.$
$10\leq n<m-$ 1:
In this case, as $t\to 0$, we have
since $j=n+1$ is the only non-zero term in the summation.
$n=-1$:
In this case, we have
$-m\leq n\leq-2$:
In this case, we have
By the identity
(A.4) $\sum_{j=0}^{l}\binom{m}{j}\frac{(n+1)_{j}}{m!}\binom{m-j}{k}(-2n-2)_{k}\\ =\begin{dcases}\frac{1}{(m-n-1)!},&\text{ when }l=-n-1\\ 0,&\text{ when }l<-n-1\end{dcases},$
where $l=j+k$ and $-m\leq n\leq-2$, we conclude that
$n=-(m+1)$:
In this case,
[…]
As $t\to 0$, we have
[…]
By the identity
(A.5) $\sum_{j=0}^{m}\binom{m}{j}\frac{(-m)_{j}(2m)_{m-j}}{m!}=1,\quad m>0,$
we conclude that, as $t\to 0$,
$n<-(m+1)$:
In this case, by (A.1)
Gathering all these computations, we obtain
$\frac{1}{m!}\frac{\partial^{m}A}{\partial t^{m}}(0,\theta)$ $=\lim_{t\to 0}\Bigg{(}-\sum_{-m-1\leq n\leq-2}\frac{1}{(m+n+1)!}\frac{\partial^{m+n+1}}{\partial t^{m+n+1}}b_{-n-2}e^{i(n+1)\theta}$
Finally, we prove the two combinatorial identities (A.4) and (A.5) used in the proof above.
Proof of (A.5).
$=\sum_{j=0}^{m}(-1)^{j}\frac{m}{m+j}\binom{m+j}{j}\binom{2m}{m+j}$
$=\sum_{j=0}^{m}(-1)^{j}\frac{m}{m+j}\binom{m}{j}\binom{2m}{m}$
$=m\binom{2m}{m}\sum_{j=0}^{m}\frac{(-1)^{j}}{m+j}\binom{m}{j}=1,$
where the last line follows from
[…]
Proof of (A.4).
For each $-n-1\leq l=j+k\leq m$,
because $(-n-2+j)_{l+n}$ is a polynomial of $j$ of degree $0\leq l+n<l$. ∎

### v4

# Singularities on maxfaces constructed by node-opening

@@ Abstract.
The node-opening technique, originally designed for constructing minimal surfaces, is adapted to construct a rich variety of new maxfaces of high genus that are embedded outside a compact set and have arbitrarily many catenoid or planar ends, thus removing the scarcity of examples of maxfaces. The surfaces look like spacelike planes connected by small necks. Among the examples are maxfaces of the Costa–Hoffman–Meeks type. Although very fruitful, the main challenge of this paper is not the construction itself, but the analysis of the positions and natures of singularities on these maxfaces. More specifically, we conclude that the singular set form curves around the waists of the necks. In generic and some symmetric cases, all but finitely many singularities are cuspidal edges, and the non-cuspidal singularities are swallowtails evenly distributed along the singular curves.
@@ 1. Introduction
Maximal surfaces are zero mean curvature immersions in the Lorentz-Minkowski space ${\mathbb{E}}_{1}^{3}$. These surfaces emerge as solutions to the variational problem of locally maximizing the area among spacelike surfaces. They share several similarities with minimal surfaces in ${\mathbb{R}}^{3}$. For instance, both are critical points of the area functional and both admit Weierstrass-Enneper representations. However, while there are rich examples of complete minimal surfaces, the only complete maximal immersion is the plane [umehara2006].
@@
Now that singularities are allowed, there are many examples of complete maxfaces, such as the Lorentzian catenoid and Kim-Yang toroidal maxface [kim2006]. In 2006, Kim and Yang [kim2006] constructed complete maximal maps of genus $k\geq 1$. When $k=1$, it is a complete embedded (in a wider sense) maxface known as Kim-Yang toroidal maxface. When $k>1$, they are not maxfaces. In [fujimori2009], the authors constructed a family of complete maxfaces $f_{k}$ for $k\geq 1$ with two ends; but when $k>1$, these may not be embedded (in a wider sense). Moreover, in 2016, Fujimori, Mohamed, and Pember [fujimori2016h] constructed maxfaces of any odd genus $g$ with two complete ends (if $g=1$, the ends are embedded) and maxfaces of genus $g=1$ and three complete embedded ends. In 2024, the third author, along with Bardhan and Biswas, proved the existence of higher genus maxfaces with one Enneper end in [BBP].
@@
To the best of our knowledge, all higher-genus maxfaces in the literature have only two or three ends, usually of the catenoid type. Very few known higher-genus maxfaces are embedded (in a wider sense). Open problems in [fujimori2009] express the hope for a large collection of examples of complete maxfaces that are embedded (in a wider sense), with higher genus and many ends.
The scarcity of examples is surprising. After all, the minimal surfaces and maxfaces admit similar Weierstrass-Enneper representations. But to construct higher genus embedded (in a wider sense) maxfaces with the required type of ends, it is usually not as direct as simply manipulating the Weierstrass data. One challenge is the singularities: While proposing the Weierstrass data, we must ensure that the singular curve does not approach the ends and is compact. Another challenge is the period problem, as illustrated in the following example.
Example 1.2 (See [BBP]*Section 1.1).
Consider the Costa minimal surface:
$M=\{(z,w)\in\mathbb{C}\times\mathbb{C}\cup\{(\infty,\infty)\}\mid w^{2}=z(z^{2}-1)\}\setminus\{(0,0),(\pm 1,0)\}$
with the data $\{\frac{a}{w},\frac{2a}{z^{2}-1}\mathrm{d}z\}$, where $a\in\mathbb{R}^{+}\setminus\{0\}$. If there exists a companion maxface of the Costa surface, then its data should be $\{-\frac{ia}{w},\frac{2ia}{z^{2}-1}\mathrm{d}z\}$ defined on $M$. Let $\tau$ be a one-sheeted loop around $(-1,0)$ that does not contain $(1,0)$. Then, $\int_{\tau}\left(\frac{2ia}{z^{2}-1}\mathrm{d}z\right)=2ia$, where $a\neq 0$. Therefore, the period problem is not solved for the corresponding maxface. So the Costa surface does not have a companion maxface.
@@
In this paper, we adapt the node-opening technique to construct a rich variety of complete maxfaces of high genus that are embedded (in a wider sense) and have an arbitrary number of spacelike ends, thus removing the scarcity of examples.
@@
The node-opening technique is a Weierstrass gluing method developed by Traizet [traizet2002e]. It constructs a family of surfaces depending on a real parameter $t$. The approach starts at $t=0$ with Weierstrass data defined on a Riemann surface with nodes, then “deforms” to Riemann surfaces for $t>0$ by opening the nodes into necks and, at the same time, “deforms” the corresponding Weierstrass data using the Implicit Function Theorem. The node-opening technique has been very successful in constructing a rich variety of minimal surfaces [traizet2002e]. To the best of our knowledge, the current paper marks the first application of the technique to surfaces in the Lorentz-Minkowski space.
@@
The Weierstrass gluing method has several advantages over other methods. On the one hand, in the existing literature on maxfaces, authors often need to assume symmetries to make the construction possible, hence only produce examples restricted to symmetries. The gluing technique has been a very powerful tool to break symmetries in minimal surfaces; in some sense, the technique was developed for this purpose [traizet2002e]. We will see later that it is equally powerful in breaking symmetries for maximal surfaces, hence ideal for removing the scarcity of examples. On the other hand, while the PDE gluing method is also popular for constructing minimal surfaces, the existence of singularities makes it difficult to be adapted for maxfaces. More specifically, one needs to identify (glue) two curves in the process, but the analysis would be difficult if the curves contain singularities. In the Weierstrass gluing process, we instead identify (glue) two annuli; hence, we can bypass the singularities.
The node-opening construction for maxfaces turns out to be very similar to that for minimal surfaces, so we will only provide a sketch. We will first give a Weierstrass data in Section 4, leaving many parameters to be determined later in Appendix A by solving the divisor problem and the period problem using the Implicit Function Theorem.
@@
This similarity implies that, for any minimal surface constructed by opening nodes, there is a corresponding maxface also constructed by opening nodes. This correspondence between maximal and minimal immersions is different from the usual correspondence through Weierstrass data [umehara2006]. In Section 3, by simply comparing notes, we obtain a rich variety of new maxfaces with high genus and arbitrarily many space-like ends, thereby remove the scarcity of examples of maxfaces. Among the examples are the Lorentzian Costa and Costa–Hoffman–Meeks (CHM) surfaces and their generalizations with arbitrarily many ends, providing positive answers to the open problems in [fujimori2009]. To the best of our knowledges, this is the first time that Lorentzian analogues of CHM surfaces were constructed.
@@
We see the the node-opening construction itself is very fruitful, but technically not that exciting. In the current paper, most effort is devoted to the more challenging task of analyzing the singularities on the constructed maxfaces.
Complete non-planar maxfaces always appear with singularities, such as cuspidal edges, swallowtails, cuspidal crosscaps, and cone-like singularities, to name a few. We refer readers to [umehara2006, sai2022, kumar2020] to explore various singularities on maxfaces. The nature of singularities can be told from the Weierstrass data. However, for the maxfaces we construct, the Weierstrass data is not given explicitly; rather, its existence is implied using the Implicit Function Theorem. Hence the analysis of the singularities must also be performed implicitly, which is a challenging task. Nevertheless, we managed to perform the analysis in various situations.
To prepare for the analysis, we first recall in Section 5.1 the governing functions whose derivatives determine the nature of singularities. Then in Section 5.2, we perform an elaborate calculation of the higher order derivatives of the governing function. It can be seen as generalizing Traizet’s calculation [traizet2008t] of the first derivative of the height differential, hence might have other use in similar node opening constructions. These allow us to conclude various results about the nature of singularities in Section 5.3, which we summarize below.
Components of the singular set are waists around the necks. In Theorem 5.1, we prove that around a specific neck and for sufficiently small, non-zero $t$, either the singular set is mapped to a single point (cone-like singularity), or all but a finite number of singular points are cuspidal edges. Moreover, the finitely many non-cuspidal singularities are generalized $A_{k}$ singularities, their positions on the waist depend analytically on $t$, and their types do not vary for sufficiently small $t$. Then, in Proposition 5.2, we show that, generically, there are four swallowtail singularities around a neck. The non-generic cases are generally hard to study, but in Section 5.3.3, we managed to analyze the singularities in the presence of rotational and reflectional symmetries. In Section 3, we will use the Lorentzian Costa and Costa–Hoffman–Meeks surfaces to exemplify our results on singularities.
@@ 2.1. Node-opening construction
We now state our first main result.
@@
The balance and non-degeneracy conditions for maxfaces turn out to be exactly the same to those for minimal surfaces, and many balanced configurations have been found when constructing minimal surfaces. So we obtain a rich variety of new maxfaces simply by comparing notes. Some of the examples are listed in Section 3. Thereby we remove the scarcity of examples of maxfaces.
Although very fruitful, the construction (proof of Theorem 2.2) is straightforward. In particular, the proof is very similar to that for minimal surfaces [traizet2002e], with only slight modifications. So we will only provide an sktech. In Section 4, we will give a Weierstrass data with undetermined parameters. In Appendix A, we will sketch the use of Implicit Function Theorem to find parameters that solve the divisor problem and the period problem.
2.2. Singularities
Let us first define
@@ 2.2. Singularities
Our main result about singularities are summarized below.
Theorem 2.3.
@@
    Assume that the configuration has rotational symmetry of order $r>1$ around a neck and $R^{(r-1)}_{l,k}\neq 0$, then for sufficiently small non-zero $t$, there are $2r$ swallowtail singularities around the neck, and they tend to be evenly distributed as $t\to 0$.
@@ 3.2. Singularities on Lorentzian Costa and CHM surfaces
4. Weierstrass data
@@
We propose the Weierstrass data in the following, as they will be useful later for the analysis of singularities. Some parameters are left undetermined. They will be determined later in Appendix A using the Implicit Function Theorem.
4.1. The Riemann Surface
@@
4.2. Gauss map and local coordinates
@@ 4.3. The height differential
$\Omega_{l}:=\{z\in{\mathbb{C}}_{l}:|v_{l,i}|\geq\varepsilon\quad\forall 1\leq i\leq n_{l}\quad\text{and}\quad|w_{l-1,j}|\geq\varepsilon\quad\forall 1\leq j\leq n_{l-1}\},$
@@
So, it suffices to prescribe the residue $-R_{l}$ and the periods around $\gamma_{l,k}$ for $1<k\leq n_{l}$.
@@
The next step of the construction is to determine the parameters using the Implicit Function Theorem. More specifically, with all parameters varying in a neighborhood of their initial values (at $t=0$), we need to prove the following
  • There exist unique values for $\alpha$ and $\beta$, depending analytically on other parameters such that the divisor conditions are satisfied.
  • With $\alpha$ and $\beta$ given above, there exist unique values for $r$, depending analytically on remaining parameters, such that the vertical period condition are satisfied.
  • With $\alpha$, $\beta$ and $r$ given above, there exist unique values for $a$, $b$ and $R$, depending smoothly on $\tau$, such that $\sum_{l}R_{l}=0$ and the horizontal period condition are satisfied.
The proofs for these claims are very similar to those for minimal surfaces [traizet2002e] with only slight modifications, and they are not essential for the following analysis of singularities, so we postpone them to Appendix A.
5. Singularities
The rest of the paper is devoted to the analysis of singularities. The study of minimal surfaces usually avoids singularities, but complete non-planar maxfaces always appear with singularities.
with $1\leq l\leq N-1,1\leq k\leq n_{l}$. We aim to analyze the nature of these singularities.
5.1. The governing function
We need to study the function
where $f_{t}:=dh_{t}/dv$.
On the other hand, it was shown in [umehara2006] that
In particular, $p=te^{i\theta}$ is a cuspidal singularity whenever
and $p$ is a swallowtail singularity whenever
5.2. Derivatives of ${\mathcal{A}}$
From equation 5.4, it is clear that we will need higher derivatives of $\mathcal{A}$ over $t$ to determine the nature of singularities. The following calculation can be seen as generalizing Traizet’s calculation [traizet2008t] of the first derivative of $dh$ over $t$.
The height differential $dh_{t}$ for the maxface, as defined in Section 4.3, has the following Laurent expansion in the annulus $t^{2}/\varepsilon<v<\varepsilon$:
where
Moreover, for each $n\in{\mathbb{Z}}$,
$a_{n}=\frac{1}{2\pi i}\int_{|v|=\varepsilon}\frac{dh_{t}}{v^{n+1}}=\frac{-1}{2\pi i}\int_{|w|=\varepsilon}\frac{dh_{t}w^{n+1}}{t^{2n+2}}$
because of the gluing. Therefore, we have
(5.5) $\frac{\partial^{m}a_{n}}{\partial t^{m}}=\frac{-1}{2\pi i}\sum_{j=0}^{m}\int_{|w|=\varepsilon}w^{n+1}\binom{m}{j}\frac{\partial^{m-j}dh_{t}}{\partial t^{m-j}}(-2n-2)_{j}\frac{1}{t^{2n+2+j}}.$
We now prove the following formula for partial derivatives of ${\mathcal{A}}$ over $t$.
Proof.
Recall the expression (5.1) of ${\mathcal{A}}$
(5.7) ${\mathcal{A}}(t,\theta)=te^{i\theta}\frac{dh_{t}}{dv}(te^{i\theta})=\sum_{n\in{\mathbb{Z}}}a_{n}(t)t^{n+1}e^{i(n+1)\theta}=\sum_{n\in{\mathbb{Z}}}b_{n}(t)t^{n+1}e^{-i(n+1)\theta},$
which can be expanded in terms of $t$ as follows
(5.8) ${\mathcal{A}}(t,\theta)=\sum_{m\geq 0}\frac{1}{m!}\frac{\partial^{m}{\mathcal{A}}}{\partial t^{m}}(0,\theta)t^{m}.$
From (5.7), we have
$n>m-1$:
$n=m-1$:
In this case, as $t\to 0$, we have
since derivatives of $a_{n}$ are bounded and, for $j<m$, $t^{m-j}\to 0.$
$10\leq n<m-$ 1:
In this case, as $t\to 0$, we have
since $j=n+1$ is the only non-zero term in the summation.
$n=-1$:
In this case, we have
$-m\leq n\leq-2$:
In this case, we have
By the identity
(5.9) $\sum_{j=0}^{l}\binom{m}{j}\frac{(n+1)_{j}}{m!}\binom{m-j}{k}(-2n-2)_{k}\\ =\begin{dcases}\frac{1}{(m-n-1)!},&\text{ when }l=-n-1\\ 0,&\text{ when }l<-n-1\end{dcases},$
where $l=j+k$ and $-m\leq n\leq-2$, we conclude that
$n=-(m+1)$:
In this case,
[…]
As $t\to 0$, we have
[…]
By the identity
(5.10) $\sum_{j=0}^{m}\binom{m}{j}\frac{(-m)_{j}(2m)_{m-j}}{m!}=1,\quad m>0,$
we conclude that, as $t\to 0$,
$n<-(m+1)$:
In this case, by (5.5)
Gathering all these computations, we obtain
$\frac{1}{m!}\frac{\partial^{m}{\mathcal{A}}}{\partial t^{m}}(0,\theta)$ $=\lim_{t\to 0}\Bigg{(}-\sum_{-m-1\leq n\leq-2}\frac{1}{(m+n+1)!}\frac{\partial^{m+n+1}}{\partial t^{m+n+1}}b_{-n-2}e^{i(n+1)\theta}$
The proof above involves two combinatorial identities, namely Eqs (5.9) and (5.10), which we now prove.
Proof of (5.10).
$=\sum_{j=0}^{m}(-1)^{j}\frac{m}{m+j}\binom{m+j}{j}\binom{2m}{m+j}$
$=\sum_{j=0}^{m}(-1)^{j}\frac{m}{m+j}\binom{m}{j}\binom{2m}{m}$
$=m\binom{2m}{m}\sum_{j=0}^{m}\frac{(-1)^{j}}{m+j}\binom{m}{j}=1,$
where the last line follows from
[…]
Proof of (5.9).
For each $-n-1\leq l=j+k\leq m$,
because $(-n-2+j)_{l+n}$ is a polynomial of $j$ of degree $0\leq l+n<l$. ∎
5.3. Nature of singularities
We are now in the possition to analyze the nature of singularities on the maxfaces we constructed by node-opening.
5.3.1. Non-cuspidal singularities
We have proved the following
Theorem 5.1.
5.3.2. Swallowtails
We first remove the trivial solutions $t=0$ by considering the function
In Section 5.2, we proved (5.6) for the partial derivatives, which we repeat below
where
In particular,
Note that, at $t=0$, we have
Then, we must look at the next derivative, namely
Proposition 5.2.
5.3.3. Symmetries
Proposition 5.3.
Proof.
Proposition 5.4.
Proof.
Remark 5.5.
Remark 5.6.
Proposition 5.7.
Proof.
Appendix A Using the Implicit Function Theorem
In Section 4, we have proposed the Weierstrass data, namely a Riemann surface $\Sigma$, Gauss map $g$, and a holomorphic 1- form $dh$, with undetermined parameters $\alpha_{l,k}$, $\beta_{l,k}$, $a_{l,k},b_{l,k},r_{l,k},R_{l},$ and $t$. In this section will prove that for sufficiently small $t$, we can find values for all the parameters, as smooth functions of $t$, such that the triplet $(\Sigma,g,dh)$ becomes a Weierstrass data for a maxface. As the argument is similar to that of [traizet2002e], we will only provide a sketch and highlight the necessary changes.
@@ Appendix A Using the Implicit Function Theorem
Proposition A.1 (Divisor condition).
@@
Proposition A.2 (Vertical periods).
@@
Proposition A.3 (Horizontal periods).
Given a balanced and rigid configuration $(p,Q)$ such that the map $Q\to W$ has rank $1$. Assume that $(\alpha,\beta,r)$ are given by previous propositions. For $\tau$ in a neighborhood of 00, there exists unique values for $a$, $b$, and $R$, depending smoothly on $\tau$, such that $\sum_{l}R_{l}=0$ and the horizontal period condition (4.2) are satisfied over the curves $\Gamma_{l,k}$ and $\gamma_{l,k}$, $1\leq l<L$ and $1<k\leq n_{l}$. Moreover, at $\tau=0$, up to a translation in ${\mathbb{C}}_{l}$, we have $a_{l,k}=\overline{b_{l,k}}=\overline{p_{l,k}}$ if $l$ is odd, $=p_{l,k}$ if $l$ is even, and $R_{l}=Q_{l}$.